2.5points

eBook

Item 3

Item 3 2.5 points Item Skipped

Dehner Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on direct labor-hours. The company based its predetermined overhead rate for the current year on the following data:

Total direct labor-hours 54,000
Total fixed manufacturing overhead cost $ 259,200
Variable manufacturing overhead per direct labor-hour $ 6.00

Recently, Job P951 was completed with the following characteristics:

Number of units in the job 25
Total direct labor-hours 100
Direct materials $ 640
Direct labor cost $ 5,400

The unit product cost for Job P951 is closest to: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Total variable manufacturing overhead estimated=(6*54,000)=$324,000

Total overhead estimated=Total variable manufacturing overhead estimated+Total fixed manufacturing overhead cost

=324,000+259,200=583200

Hence predetermined overhead rate=Total overhead estimated/Total direct labor-hours estimated

=583200/54,000=$10.8/direct labor hour

Hence applied overhead=(10.8*100)=$1080

Add:Direct materials=640

Add:Direct labor cost=5400

Total cost=$7120

Hence unit product cost =7120/25

=$284.8 per unit